OpenStudy (anonymous):

Rewrite the expression using radical notation 272/3

ganeshie8 (ganeshie8):

\(\huge 27^{\frac{2}{3}}\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

YEA WHATS THE ASNWER

ganeshie8 (ganeshie8):

if so, use below property : \(\huge \mathbb{ a^{\frac{m}n{}} = \sqrt[n]{a^m}}\)

OpenStudy (anonymous):

I NEW TO THIS CLASS I NEED A ANSWERS QUICK

ganeshie8 (ganeshie8):

are you trying to simplify the expression, or just want to rewrite it into radical form ?